Suppose market demand is given Qd=a-bp as and market supply is given as, Qs=c+dp and then find the following

o  Equilibrium price

o  Equilibrium quantity

o  Price elasticity of demand at equilibrium

o  Price elasticity of supply at equilibrium 


1
Expert's answer
2021-11-24T12:38:31-0500

Price elasticity of supply at equilibrium


"price elasticity of demand=\frac{percentage change in quantity supplied}{percentagechange in price}"


PERCENTAGE CHANGE IN QUANTINTY SUPPLIED ="\frac {\Delta Q}{Q}"

PERCENTAGE CHANGE IN PRICE="\frac {\Delta P}{P}"

change in quantity SUPPLIED "\Delta Q =Q{1}- Q"

change in price"\Delta Q =P{1}- P"


PRICE ELASTICITY ="\frac{\Delta Q}{ Q} \times \frac{P}{\Delta P}"


PRICE ELASTICITY ="\frac{\Delta Q}{\Delta P } \times \frac{P}{Q}" 

but from the supply equation "\frac{\Delta Q}{\Delta P }" is the gradient of the equation


therefore from "Q{s}=(c+ dp)" 

"d" is the gradient of the equation



the price elasticity of supply will be


"P{e}=d \times \frac{P}{q}"